6. Shireen measures the gauge pressure, pa 3 atm, volume, V 2 m and temperature T 126.85 °C, of an ideal gas. What is the correct number of moles of this gas in terms of the gas constant R (assume 1 atm 1 x 105 Pa; 1 Pa 1 N/m2) A. 500 N.m.K R B. 1000 K-1 N.m.K 1500 N.ma.K D. 2000 N
Answer #1

From the ideal gas equation, we know that

PV = nRT\Rightarrow n = \frac{PV}{RT}=\frac{(3+1)*1.01*10^5*2}{R*(273.15+126.85)}= \frac{2000}{R}N-m/K